Molality of a solution having 14.2 gram Na2Sodissolved in 2 kg solution

Grade:11

1 Answers

Vikas TU
4 years ago
The molality of a solution – is the number of moles of solute in 1 kg of solvent (denoted mol/kg).
First we have to calculate the molar mass of Na2SO4 (M):
M = 23.0*2 + 32.1 + 16.0*4 = 142.1.
Therefore 14.2 g of it contain 14.2 / 142.1 = 0.099 mol.
So we have 0.099 mol of Na2SO4 in 2 kg of water. 
Than 1 kg of water contains 0.09mol / 2kg = 0.0499 mol/kg